OpenStudy (anonymous):

If i have a question like 3(5^x+1) = 15 I ended up with 15^x-1 = 15 If I then drop the bases since they are equal does the 15 on the rights exponent end up as 1?

OpenStudy (anonymous):

is your question is\[3(5^x+1)=15\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you should proceed like 3(5^x+1)=15 5^x+1=15/3; 5^x+1=5 5^x=4 \[x=\log_{5}{4}\]

OpenStudy (anonymous):

i don't know how did u get 15^x-1=15???

OpenStudy (anonymous):

the answer in the book is 0. I wrote that wrong it was still a +.

OpenStudy (anonymous):

if u need zero then the question should look like:\[3*5^{(x+1)}=15\]

OpenStudy (anonymous):

check it...otherwise you can't get 0

OpenStudy (anonymous):

Yes, im sorry, how did you write that?

OpenStudy (anonymous):

\[3*5^{(x+1)}=15:\]\[5^{x+1}=\frac{15}{3}\]\[5^{x+1}=5\]\[x+1=1;\]x=0

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I couldn't just multiply the 3 with the 5 to get 15?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

its not 5 its 5^(x+1)

OpenStudy (anonymous):

for example \[3*5^2\neq (3*5)^2\]

OpenStudy (anonymous):

Order of operations, you must deal with the exponent before you can start multiplying.

OpenStudy (anonymous):

i think you need a little study on the properties of indices...

OpenStudy (anonymous):

I do. 10 years from math and i've lost most of these things. im trying to get th em back through trial and error, since I dont have time to redo earlier classes.

OpenStudy (anonymous):

okay..carry on...don't give up...someone is always here to help you.

OpenStudy (anonymous):

I appreciate it greatly!

OpenStudy (anonymous):

If i may, dipankarstudy just made x a value so that the 5^(x+1) would equal 5, and thus it would be easy to manipulate the 3*5. If you can see he was able to quickly solve for x because of that, it makes the problem much easier.